# A Non-Relativistic Limit for Heterotic Supergravity and its Gauge Lagrangian

### v3

@@ 4.2 Gauged terms
Before analyzing the gauge contributions of the heteorotic Lagrangian, let’s focus on the ungauged terms depending on the b-field,
$L_{b}$ $=$ $-\frac{1}{4}\partial_{\mu}{b_{\nu\rho}}\partial_{\sigma}{b_{\gamma\epsilon}}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}+\epsilon_{ab}\eta^{ac}\partial_{\mu}{\tau_{\nu}{}^{b}}\partial_{\rho}{b_{\sigma\gamma}}\tau^{\sigma}{}_{c}h^{\mu\rho}h^{\nu\gamma}+\frac{1}{2}\partial_{\mu}{b_{\nu\rho}}\partial_{\sigma}{b_{\gamma\epsilon}}h^{\mu\gamma}h^{\nu\sigma}h^{\rho\epsilon}$ (4.6)
$-\epsilon_{ab}\eta^{ac}\partial_{\mu}{\tau_{\nu}{}^{b}}\partial_{\rho}{b_{\sigma\gamma}}\tau^{\rho}{}_{c}h^{\mu\sigma}h^{\nu\gamma}-\epsilon_{ab}\eta^{ac}\partial_{\mu}{\tau_{\nu}{}^{b}}\partial_{\rho}{b_{\sigma\gamma}}\tau^{\sigma}{}_{c}h^{\mu\gamma}h^{\nu\rho}\,.$
These terms are not gauge invariant, due to the transformation (3.11). Therefore, we rewrite these contributions in terms of the new curvature for the b-field,
$-\frac{1}{12}\bar{h}_{\mu\nu\rho}\bar{h}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}{}^{b}\bar{h}_{\rho\sigma\gamma}\tau^{\sigma}{}_{c}h^{\mu\gamma}h^{\nu\rho}+\nabla_{\mu}\tau_{\nu}^{-}f_{\gamma\rho}{}^{i}\alpha_{-i}h^{\mu\gamma}h^{\nu\rho}+L_{\textrm{comp}}\,,$ (4.7)
where $L_{\textrm{comp}}=L^{(2)}_{\textrm{non-cov}}+L^{(3)}_{\textrm{non-cov}}+L^{(4)}_{\textrm{non-cov}}$ with
$L^{(2)}_{\textrm{non-cov}}$ $=$ $\nabla_{\mu}\tau_{\nu}^{-}a_{[\rho}^{i}\nabla_{\gamma]}\alpha_{-i}h^{\mu\gamma}h^{\nu\rho}+3\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}a_{[\sigma}^{i}\alpha_{-i}\nabla_{\rho}\tau_{\gamma]}^{-}\tau^{\sigma}_{c}h^{\mu\gamma}h^{\nu\rho}$ (4.8)
$+\frac{1}{2}h_{\mu\nu\rho}a_{\gamma}^{i}\alpha_{-i}\nabla_{\sigma}(\tau_{\epsilon}^{-})h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}\,,$
$L^{(3)}_{\textrm{non-cov}}=-\nabla_{\mu}\tau_{\nu}^{-}f^{i}{}_{jk}a_{\gamma}^{j}a_{\rho}{}^{k}\alpha_{-i}h^{\mu\gamma}h^{\nu\rho}\,,$ (4.9)
and
$L^{(4)}_{\textrm{non-cov}}$ $=$ $\frac{3}{4}a_{[\nu}^{i}\alpha_{-i}\nabla_{\mu}(\tau_{\rho]^{-}})a_{\gamma}^{j}\alpha_{-j}\nabla_{\sigma}\tau_{\epsilon}^{-}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}\,.$ (4.10)
In the next subsection we will study the contributions to the gauge Lagrangian coming from the Yang-Mills terms and the Chern-Simons contributions. We split the presentation according to an expansion in gauge fields.
4.3 Quadratic gauged contributions
4.3.1 Quadratic Yang-Mills contributions
The quadratic contributions coming from $-\frac{1}{4}\hat{F}^{2}$ are given by
$-\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-\alpha_{-i}\alpha_{-}^{i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\mu}_{a}\tau^{\rho}_{b}h^{\nu\sigma}$
$+2\alpha_{-i}\eta^{ab}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau_{\sigma}^{-}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-2\alpha_{-i}\nabla_{\mu}\tau_{\nu}{}^{-}\nabla_{[\rho}a_{\sigma]}^{i}h^{\mu\rho}h^{\nu\sigma}\,.$ (4.11)
The previous expression can be written in a covariant form as
$-\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-\alpha_{-i}\alpha_{-}^{i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\mu}_{a}\tau^{\rho}_{b}h^{\nu\sigma}$
$-2\alpha_{-i}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau^{\nu}_{+}h^{\mu\rho}-\alpha_{-i}\nabla_{\mu}\tau_{\nu}{}^{-}f_{\rho\sigma}^{i}h^{\mu\rho}h^{\nu\sigma}+L^{(3)}(a,\alpha)$ (4.12)
where the inclusion of the (gauge) covariant derivative and the curvature $f_{\mu\nu i}$ introduces cubic contributions
$L^{(3)}(a,\alpha)=-\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}f^{i}{}_{jk}a^{j}_{\rho}a^{k}_{\sigma}h^{\mu\rho}h^{\nu\sigma}+2\alpha_{-i}\eta^{ab}f^{i}{}_{jk}a_{\mu}^{j}\alpha_{-}^{k}\nabla_{[\nu}\tau_{\rho]}^{-}\tau_{\sigma}^{-}\tau_{a}^{\nu}\tau_{b}^{\sigma}h^{\mu\rho}\,.$ (4.13)
4.3.2 Quadratic Chern-Simons contributions
The quadratic contributions coming from $-\frac{1}{12}\bar{H}^{2}$ are given by
$-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\mu}_{+}\tau^{[\sigma}_{c}h^{\rho]\nu}-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}^{b}}\nabla_{\rho}{\tau_{\sigma}^{-}}\tau^{\nu}_{+}\tau^{[\rho}_{c}h^{\sigma]\mu}$
$+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\sigma}_{b}\tau^{[\nu}_{a}h^{\mu]\rho}+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{[\mu}_{b}h^{\nu]\sigma}$
$-\alpha_{-}^{i}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}a_{\sigma]i}h^{\mu\sigma}h^{\nu\rho}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}h_{\rho\sigma\gamma}\tau^{\rho}_{+}h^{\mu\sigma}h^{\nu\gamma}-L^{(2)}_{\textrm{non-cov}}\,.$ (4.14)
Among these terms we have identified all the quadratic non-covariant contributions needed in (4.7). We can write the previous expression in covariant form as
$-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\mu}_{+}\tau^{[\sigma}_{c}h^{\rho]\nu}-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}^{b}}\nabla_{\rho}{\tau_{\sigma}^{-}}\tau^{\nu}_{+}\tau^{[\rho}_{c}h^{\sigma]\mu}$
$+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\sigma}_{b}\tau^{[\nu}_{a}h^{\mu]\rho}+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{[\mu}_{b}h^{\nu]\sigma}$
$-\alpha_{-}^{i}\nabla_{\mu}{\tau_{\nu}^{-}}f_{\rho\sigma i}h^{\mu\sigma}h^{\nu\rho}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}\bar{h}_{\rho\sigma\gamma}\tau^{\rho}_{+}h^{\mu\sigma}h^{\nu\gamma}-L^{(2)}_{\textrm{non-cov}}+\tilde{L}^{(3,4)}(a,\alpha)\,,$ (4.15)
where once again the inclusion of the curvatures $f_{\mu\nu i}$ and $\bar{h}_{\rho\sigma\gamma}$ in the last line of the previous expression introduces cubic and quartic contributions, respectively, that we will take into account when we analyze the next orders. We capture all these new contributions under the notation $\tilde{L}^{(3,4)}(a,\alpha)$, which splits in the following form
$\tilde{L}^{(3)}(a,\alpha)=-\alpha_{-}^{i}\nabla_{\mu}\tau_{\nu}^{-}f_{i}{}^{jk}a_{\rho j}a_{\sigma k}h^{\mu\sigma}h^{\nu\rho}\,,$ (4.16)
and
$\tilde{L}^{(4)}(a,\alpha)=\frac{3}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}a_{[\sigma}^{j}\alpha_{-j}\nabla_{\rho}(\tau_{\gamma]}^{-})\tau_{+}^{\rho}h^{\mu\sigma}h^{\nu\gamma}\,.$ (4.17)
4.4 Cubic gauge contributions
4.4.1 Cubic Yang-Mills contributions
The cubic contributions coming from $-\frac{1}{4}\hat{F}^{2}$ are given by
$-L^{(3)}_{\textrm{non-cov}}-L^{(3)}(a,\alpha)-\tilde{L}^{(3)}(a,\alpha)\,.$ (4.18)
The previous expression exactly cancels the contributions obtained in (4.13) and (4.16) and there are no additional contributions at this order in fields.
4.5 Quartic gauge contributions
4.5.1 Quartic Chern-Simons contributions
The quartic terms in $-\frac{1}{12}\bar{H}^{2}$ are given by
$-\tilde{L}^{(4)}(a,\alpha)\,,$ (4.19)
and there are no other additional contributions at this order, or higher, in fields.
@@ 5 Discussion
In this work we have studied the finite form of the heterotic supergravity Lagrangian, extending the formulation of [13]. The Ricci scalar part of the Lagrangian has been rewritten in terms of $\nabla_{\mu}\tau_{\nu}{}^{a}$, $\nabla_{\mu}\tau^{\nu}{}_{a}$, $Q_{\mu\nu}{}^{a}$ (eq. (3.4)) and $S_{\mu}{}^{\nu}{}_{a}$ (eq. (3.6)), showing the explicit diffeomorphism invariance of the theory.
When we include the heteoric gauge field, the bosonic part of the Lagrangian of this theory contains two divergent parts given by the Chern-Simons terms and the Yang-Mills Lagrangian which exactly compensate each other, as it happens in bosonic supergravity, giving rise to a finite Lagrangian. The gauge Lagrangian depends on a pair of gauge fields, $a_{\mu}^{i}$ and $\alpha_{-}^{i}$, the former transforming as a gauge connection, so we constructed its curvature, and the latter transforming as a gauge vector, so we do not need to construct a curvature. Interestingly enough, the Green-Schwarz mechanism for the b-field can be trivialized in this limit and therefore one can define $\tilde{b}_{\mu\nu}$ as in (3.16), and its curvature $\tilde{h}_{\mu\nu\rho}$, both of them gauge invariant. The gauge contributions in $\tilde{h}_{\mu\nu\rho}$ can be though of as Chern-Simons contributions.
In this work we have presented the finite Lagrangian of the theory which contains quadratic, cubic and quartic contributions. The full gauge Lagrangian is given by
$L^{(g)}_{\textrm{het}}$ $=$ $-\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-\alpha_{-i}\alpha_{-}^{i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\mu}_{a}\tau^{\rho}_{b}h^{\nu\sigma}$ (5.1)
$-2\alpha_{-i}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau^{\nu}_{+}h^{\mu\rho}-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\mu}_{+}\tau^{[\sigma}_{c}h^{\rho]\nu}$
$-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}^{b}}\nabla_{\rho}{\tau_{\sigma}^{-}}\tau^{\nu}_{+}\tau^{[\rho}_{c}h^{\sigma]\mu}+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\sigma}_{b}\tau^{[\nu}_{a}h^{\mu]\rho}$
$+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{[\mu}_{b}h^{\nu]\sigma}-\alpha_{-}^{i}\nabla_{\mu}{\tau_{\nu}^{-}}f_{\rho\sigma i}h^{\mu\sigma}h^{\nu\rho}$
$-\frac{1}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}\bar{h}_{\rho\sigma\gamma}\tau^{\rho}_{+}h^{\mu\sigma}h^{\nu\gamma}\,,$
while the ungauged action must include the Chern-Simons terms in the curvature of the b-field as we show in (4.7), i.e.,
$-\frac{1}{12}\bar{h}_{\mu\nu\rho}\bar{h}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}{}^{b}\bar{h}_{\rho\sigma\gamma}\tau^{\sigma}{}_{c}h^{\mu\gamma}h^{\nu\rho}\,.$ (5.2)

### v5

@@ 4.2 Gauged terms
Now we compute the gauge contributions of the heterotic Lagrangian given by the inclusion of the $\hat{A}_{\mu}{}^{i}$ field.
[…]
with
and $\hat{f}_{ijk}=cf_{ijk}$. In this case, the Lagrangian also remains finite after considering the limit $c\rightarrow\infty$ and it can be written as
$L=ee^{-2\varphi}\Big(R^{(0)}+4\partial_{\mu}\varphi\partial_{\nu}\varphi h^{\mu\nu}+T_{H^{2}}+T_{F^{2}}\Big)$ (4.11)
where
$T_{H^{2}}=\frac{1}{2}\alpha_{-}^{2}\eta^{ab}\nabla_{[\mu}{\tau_{\nu]}^{-}}\bar{h}_{\rho\sigma\gamma}\tau_{\epsilon}^{-}\tau^{\rho}{}_{a}\tau^{\epsilon}{}_{b}h^{\mu\sigma}h^{\nu\gamma}+2\alpha_{-}^{2}\epsilon_{ab}\eta^{ac}\eta^{de}\nabla_{[\mu}{\tau_{\nu]}^{b}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau_{\gamma}^{-}\tau^{\mu}_{d}\tau^{\sigma}_{c}\tau^{\gamma}_{e}h^{\nu\rho}$
$-\frac{1}{2}\alpha_{-}^{4}\eta^{ab}\eta^{cd}\nabla_{[\mu}{\tau_{\nu]}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau_{\gamma}^{-}\tau_{\epsilon}^{-}\tau^{\mu}_{a}\tau^{\sigma}_{c}\tau^{\gamma}_{d}\tau^{\epsilon}_{b}h^{\nu\rho}$
$+2\eta_{ab}\eta^{cd}\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{[\rho}{\tau_{\sigma]}{}^{b}}\tau^{\mu}{}_{c}\tau^{\sigma}{}_{d}h^{\nu\rho}+2\epsilon_{ab}\epsilon_{cd}\eta^{ae}\eta^{cf}\nabla_{[\mu}{\tau_{\nu]}{}^{b}}\nabla_{[\rho}{\tau_{\sigma]}{}^{d}}\tau^{\nu}{}_{f}\tau^{\sigma}{}_{e}h^{\mu\rho}$
$-\frac{1}{12}\bar{h}_{\mu\nu\rho}\bar{h}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}{}^{b}}\tau^{\rho}{}_{c}\bar{h}_{\rho\sigma\gamma}h^{\mu\sigma}h^{\nu\gamma}$ (4.12)
is the finite contribution coming from the $\bar{H}^{2}$-term, with $\alpha_{-}^{2}=\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}$, and the gauge-invariant 3-form is given by
$\bar{h}_{\mu\nu\rho}=h_{\mu\nu\rho}+3\partial_{[\mu}(a_{\nu}^{i}\tau_{\rho]}^{-}\alpha_{-i})$ (4.13)
The finite contribution coming from the $\hat{F}^{2}$-term is given by
$T_{F^{2}}$ $=$ $-2\alpha_{-}^{2}\eta^{ab}\nabla_{[\mu}{\tau_{\nu]}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-2\alpha_{-}^{i}\eta^{ab}\eta_{ij}\nabla_{\mu}{\alpha_{-}^{j}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{\sigma}_{b}h^{\mu\nu}$ (4.14)
$-f_{\mu\nu}^{i}\alpha_{-}^{j}\eta_{ij}\nabla_{\rho}{\tau_{\sigma}^{-}}h^{\mu\rho}h^{\nu\sigma}-\frac{1}{2}\eta^{ab}\eta_{ij}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{\nu}{\alpha_{-}^{j}}\tau_{\rho}^{-}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{\sigma}_{b}h^{\mu\nu}\,$,
As the reader can easily verify, all the gauge contributions to the heterotic Lagrangian in the NR limit depend on the vector $\alpha_{-}{}^{i}$.
4.3 Comparison to the Bergshoeff-Romano formulation
As an alternative to the expansion for the gauge field discussed in this paper, it is possible to construct a finite formulation by considering the following expansion [21],
$V_{\nu}{}^{i}=c^{2}\tilde{\tau}_{\nu}{}^{-}v_{-}{}^{i}+\tilde{\tau}_{\nu}{}^{+}v_{+}{}^{i}+\tilde{e}_{\nu}{}^{a^{\prime}}v_{a^{\prime}}{}^{i}\,$, (4.15)
where $V_{\mu}{}^{i}$ plays the role of the $A_{\mu}{}^{i}$ vector. In this formalism, the c-expansion of the vielbein in this formulation includes gauge field contributions and it is given by
$E_{\mu}{}^{-}$ $=$ $c\tilde{\tau}_{\mu}{}^{-},$ (4.16)
$\hat{E}_{\mu}{}^{+}$ $=$ $-c^{3}\frac{v_{-}{}^{i}v_{-i}}{2}\tilde{\tau}_{\mu}{}^{-}+c\tilde{\tau}_{\mu}{}^{+}\,$, (4.17)
$E_{\mu}{}^{a^{\prime}}$ $=$ $\tilde{e}_{\mu}{}^{a^{\prime}}\,$, (4.18)
while the expansion for the B-field is given by
$B_{\mu\nu}=c^{2}\tilde{\tau}_{\mu}{}^{a}{\tilde{\tau}_{\nu}}^{b}\epsilon_{ab}+2c^{2}\tilde{\tau}_{[\mu}{}^{-}\tilde{e}_{\nu]}{}^{a^{\prime}}v_{-}{}^{i}v_{a^{\prime}i}+\tilde{b}_{\mu\nu}\,.$ (4.19)
The gauge transformations of the gauge degrees of freedom are given by
$\delta_{\lambda}v_{+}{}^{i}$ $=$ $\partial_{+}\lambda^{i}+\sqrt{2}gf_{jk}{}^{i}\lambda^{j}v_{+}{}^{k}\,$, (4.20)
$\delta_{\lambda}v_{-}{}^{i}$ $=$ $\sqrt{2}gf_{jk}{}^{i}\lambda^{j}v_{-}{}^{k}\,$, (4.21)
$\delta_{\lambda}v_{a^{\prime}}{}^{i}$ $=$ $\partial_{a}\lambda^{i}+\sqrt{2}gf_{jk}{}^{i}\lambda^{j}v_{a^{\prime}}{}^{k}\,$. (4.22)
We can read from the previous transformations that $v_{+}{}^{i}$ and $v_{a}{}^{i}$ play the role of connections, while $v_{-}{}^{i}$ is a gauge vector.
In this prescription $\tilde{\tau}_{-}{}^{a}$ is gauge invariant but $\tilde{\tau}_{+}{}^{a}$ transforms in a non-covariant way under gauge transformations,
$\delta\tilde{\tau}_{\mu}{}^{+}=\frac{v_{-}{}^{i}\partial_{-}\lambda^{i}}{1+v_{+}{}^{i}v_{-i}}\tilde{\tau}_{\mu}{}^{-}\,.$ (4.23)
Since $v_{-}{}^{i}$ is boost-invariant and transforms as a vector under gauge transformations, we will work with the case $v_{-}{}^{i}=0$, so that all
$\tilde{\tau}_{\mu}{}^{a}=\tau_{\mu}{}^{a}\quad\quad\tilde{e}_{\mu}{}^{a^{\prime}}=e_{\mu}{}^{a^{\prime}}\,$. (4.24)
At this point, it is important to observe that taking $\alpha_{-}{}^{i}=0$ in the prescription used in the previous sections is not equivalent to $v_{-}{}^{i}=0$. In the formalism [21] this condition still produces a non-trivial Green-Schwarz mechanism for the b-field, given by
$\delta_{\lambda}b_{\mu\nu}$ $=$ $2\tau_{[\mu}{}^{+}e_{\nu]}{}^{a^{\prime}}(v_{+}{}^{i}\partial_{a^{\prime}}\lambda_{i}-v_{a^{\prime}i}\partial_{+}\lambda^{i})-2e_{[\mu}{}^{a^{\prime}}e_{\nu]}{}^{b^{\prime}}v_{a^{\prime}}{}^{i}\partial_{b^{\prime}}\lambda_{i}$ (4.25)
$-4(\tau_{[\mu}{}^{-}\tau_{\nu]}{}^{+}v_{+i}+\tau_{[\mu}{}^{-}e_{\nu]}{}^{a^{\prime}}v_{a^{\prime}i})\partial_{-}\lambda^{i}\,.$
Moreover, in the case $v_{-}{}^{i}=0$, the transformation (4.25) is not ambiguous and it cannot be eliminated with a field redefinition. When $v_{-}{}^{i}\neq 0$, one can try to use the anomalous transformation of $\tilde{\tau}_{\mu}{}^{+}$ to trivialize the Green-Schwarz mechanism.
The explicit finite gauged action in this formulation, for the case $v_{-}{}^{i}=0$ is written in the Appendix B.
@@ 5 Discussion
In this work, we study the finite form of the heterotic supergravity Lagrangian, extending the formulation of [13], and providing an alternative way to that of [21] of taking this limit. The bosonic part of the Lagrangian of this theory contains two divergent parts given by the Chern-Simons terms and the Yang-Mills Lagrangian which exactly compensate each other, as it happens in bosonic supergravity, giving rise to a finite Lagrangian. The gauge Lagrangian depends on a pair of gauge fields, $a_{\mu}^{i}$ and $\alpha_{-}^{i}$, the former transforming as a gauge connection, so we construct its curvature, and the latter transforms as a gauge vector, so we do not need to construct a curvature. While the formulation presented here is not the first non-relativistic heterotic supergravity in the literature, it is very likely that all these approaches are connected by field redefinitions. In particular, in this work we also compute part of the finite Lagrangian of the formulation given in [21] for a particular case where the vielbein is gauge invariant (Appendix B). While there is not yet a proof, a possible way to connect these two formulations of non-relativistic heterotic supergravity is to trivialize the gauge transformation of the vielbein in both theories (without imposing extra constraints), which will need field redefinitions on the b-field and the gauge fields, leading to the same Green-Schwarz mechanism.
A remarkable point of the construction here presented is that the Green-Schwarz mechanism for the b-field can be trivialized when one takes the NR limit and therefore it is possible to redefine the b-field to the new $\tilde{b}_{\mu\nu}$ as in (3.17), and its curvature $\bar{h}_{\mu\nu\rho}$, both of them are gauge invariant. While there is not yet proof, this cancellation of the Green-Schwarz transformation could indicate that the gauge anomaly could be trivially resolved under the NR limit. We will discuss this point in the outlook, as a promising future direction of work.